### v1

# Optomechanical self-organization in a mesoscopic atom array

Abstract
A fundamental challenge in modern physics is bridging the gap between an exact description of few-body microscopic systems and the emergent description of many-body macroscopic systems. This gap may be bridged by tracing the properties of well-controlled intermediate-size mesoscopic systems. Here, we study mesoscopic signatures of a spatial self-organization phase transition in deterministically prepared arrays of between 10 and 22 atoms inside an optical cavity. Through precise engineering of the atom-cavity interactions, we reveal how critical behavior depends on atom number. We identify characteristic dynamical behaviors related to symmetry breaking and system size in the self-organized regime, and observe a finite optomechanical susceptibility at the critical point. This work opens the door to probing particle-number- and time-resolved properties of phase transitions in mesoscopic systems.
The scale of a physical system plays a key role in determining its material properties. With increasing size comes increasing complexity, and complexity leads to the emergence of macroscopic phenomena that, often, cannot be extrapolated from microscopic behaviors [1]. A well-known example of this paradigm is the appearance of phase transitions in the thermodynamic limit, i.e. as a system’s particle number tends to infinity. In between the microscopic and macroscopic extremes lies the mesoscopic regime. Mesoscopic systems not only exhibit characteristics of macroscopic systems, but also host a variety of rich properties themselves. In systems of intermediate size, assumptions of local equilibrium break down and internal microstructures may affect properties on longer length scales [2]. The fact that microscopic fluctuations have a measurable influence on mesoscopic scales also makes mesoscopic systems an interesting playground for the study of critical phenomena and phase transitions. Predicted behaviors of mesoscopic systems near critical points include the smoothing out of the slope of the order parameter [3], finite-size scaling exponents [4], and switching dynamics between symmetry-broken phases [5, 6, 7].
Technical advances in engineering, probing and measuring mesoscopic systems have propelled experiments in a variety of physical platforms. Notable examples include solid state systems [8, 9], nanostructured [10] and nanomechanical systems [11, 12, 13, 14], nanomagnets [15, 16], and cold atoms and ions [17, 18, 19, 20, 21, 22, 23]. Breakthroughs in the control of atoms trapped in optical tweezer arrays represent a leap forward in precisely constructing mesoscopic, ultracold atomic systems. We show that deterministic control over particle number and interaction strength can be achieved with an optical tweezer array inside a standing-wave optical cavity. The unprecendented versatility in the positioning of individual atoms offered by tweezer arrays has enabled numerous advances in quantum simulation, sensing, and information processing [24, 25]. We use the tweezers to adjust the position of each atom with respect to the spatial structure of the cavity field, thereby precisely programming the atom-cavity coupling strength. This system uniquely lends itself to the detailed study of optomechanics in mesoscopic samples.
We focus on the mechanical self-organization generated by light-mediated forces among optically excited atoms within an optical cavity [26, 27]. When such atoms are driven by coherent light above a certain intensity, they exhibit a phase transition in which they spontaneously scatter light into the cavity. That light creates a spatially dependent potential on the atoms that causes them to self-organize and superradiantly emit into the cavity with a phase that breaks $\mathbb{Z}_{2}$ symmetry. This self-organization phase transition maps onto the Dicke model [28, 29]. Previous experiments on the optomechanical Dicke phase transition [29] and other phenomena arising from cavity-mediated forces among atoms [30, 31, 32, 33, 34] were conducted on bulk macroscopic samples, with particle numbers $N\gtrsim 10^{5}$. Here, we realize the mesoscopic limit of the Dicke phase transition in deterministically prepared tweezer arrays containing between 10 and 22 atoms. We observe a critical point that depends on atom number, a finite switching time scale between self-organized states, and a collectively altered susceptibility to external forces. This setup allows us to study the crossover from the few- to many-body regime of phase transitions, advancing our knowledge of finite size effects and dynamical behavior in mesoscopic systems.
I Cavity-Mediated interactions and Self-organization
[…]
To realize mesoscopic self-organization, we use the experimental setup schematically shown in Fig. 1A. It contains a one-dimensional tweezer array into which we load a deterministic number $N$ of 87Rb atoms. The array is aligned along the axis of a high-finesse Fabry-Pérot optical cavity (the $z$-axis). The cavity is placed in-vacuum and has a TEM00 mode whose frequency, $\omega_{\mathrm{c}}$, lies near the resonance frequency, $\omega_{\mathrm{a}}$, of the 87Rb $F=2\rightarrow F^{\prime}=3$ $D_{2}$ transition (wavelength $\lambda=780\,$nm) [35, 36]. The atoms, prepared in the $F=2$ ground state manifold, are exposed to a circularly polarized standing wave pump beam of frequency $\omega_{\mathrm{p}}$ (detuned from the atoms by $\mathrm{\Delta_{pa}}=\omega_{\mathrm{p}}-\omega_{\mathrm{a}}$ and from the cavity by $\mathrm{\Delta_{pc}}=\omega_{\mathrm{p}}-\omega_{\mathrm{c}}$) propagating along the $x$-axis. The circularly polarized pump creates a fictitious magnetic field along $x$ through vector light shifts and defines the atomic quantization axis. The pump drives the $\sigma^{-}$ electric dipole transition of the atoms and its power is linearly ramped up to the desired value in $50\,\mathrm{\mu s}$. The initial ramp optically pumps the atoms into the |$\ket{F=2,m_{F}=-2}$ stretched state, whence the atoms are excited on the cycling transition, thereby restricting spin dynamics to a two-level system. We achieve a maximal vacuum Rabi coupling of $g_{0}=2\pi\times 3.1/\sqrt{2}$ MHz between the $y$-polarized cavity field and the atomic $\sigma^{-}$ transition. The cavity half-linewidth is $\kappa=2\pi\times 0.53\,\mathrm{MHz}$, resulting in a sub-$\mu$s cavity response time defined by $1/\kappa$. Given that the atomic half-linewidth is $\gamma=2\pi\times 3.0\,\mathrm{MHz}$, we achieve a single-atom cooperativity $C=g_{0}^{2}/(2\kappa\gamma)=1.5$ [37]. Because the atom-cavity interaction depends strongly on the spatial structure of the pump and cavity standing waves [38, 39, 40, 41], the tweezers are aligned to be all within the same pump antinode wavefront (so the atoms all experience the same pump Rabi frequency, $\mathrm{\Omega}$) and are positioned at cavity nodes spaced $4.5\lambda$ apart. By spacing the tweezers by a half-integer number of wavelengths, we eliminate biasing in our system due to center-of-mass drift of the tweezer array with respect to the cavity standing wave.
Light inside the cavity creates a potential energy landscape for the atoms through the exchange of cavity photons [27]. For atom $n$ at position $\hat{z}_{n}$, the cavity interaction potential, produced by interference between the pump and cavity fields, is proportional to $S\sin(k\hat{z}_{n})(\hat{c}+\hat{c}^{\dagger})$, where $\hat{c}$ ($\hat{c}^{\dagger}$) is the bosonic annihilation (creation) operator for a cavity photon, $S\simeq\mathrm{\Omega}g_{0}/(2\mathrm{\Delta_{pa}})$ is the coherent scattering rate of a pump photon into the cavity, and $k=2\pi/\lambda$. The strength of the potential thus depends on the atom’s position inside the cavity and on the magnitude of the cavity field. When the pump is sufficiently detuned from the atomic transition and red-detuned from cavity resonance, the cavity field follows the atomic motion such that $\hat{c}+\hat{c}^{\dagger}\propto-S\sum_{m=1}^{N}\sin(k\hat{z}_{m})$. In this adiabatic regime, the effective potential is minimized when $\langle\sin(k\hat{z}_{n})\rangle=\pm 1$ for all $n$. That is, the most favorable configurations of the system are the ones where the atoms self-organize by collectively moving to cavity-field antinodes with the same sign.
The steady state of the optomechanical system is governed by competition between the cavity-mediated interaction and the tweezer potential, which, for tweezers centered at the cavity nodes, is minimized when $\langle\sin(kz_{n})\rangle=0$ for all $n$. This competition gives rise to the emergence of self-organization as a function of $S$, which we tune through $\mathrm{\Omega}$, thus defining a critical pump strength $\mathrm{\Omega_{c}}$. For $\mathrm{\Omega}<\mathrm{\Omega_{c}}$, the tweezers dominate over the cavity field and the atoms remain centered at the cavity nodes (as shown in Fig. 1A(i)). Although the atom array is uncoupled to the cavity when positioned exactly at the nodes, zero-point and thermal fluctuations shift the atom positions, causing the array to emit light incoherently into the cavity. For $\mathrm{\Omega}>\mathrm{\Omega_{c}}$, the negative curvature of the cavity potential overcomes the positive tweezer trap curvature and it becomes energetically favorable for all atoms to move towards antinodes with the same sign and for their emission to interfere constructively. The superradiant field’s phase is determined by the sign of the antinodes chosen by the atoms (Fig. 1A(ii-iii)). For atoms spaced by a half-integer number of wavelengths, adjacent atoms move in opposite directions to achieve self-organization (antiphase motion), in contrast to integer spacing where the atoms move in the same direction (center-of-mass motion). We refer to the collective mode that is linearly coupled to the cavity as the dominant mode.
To confirm the self-organization experimentally, we detect the cavity field $\hat{c}$ and observe $\mathbb{Z}_{2}$ symmetry breaking. Assuming $\mathrm{\Delta_{pa}}<0$ (the case for all data shown), light scattered by the atoms affects one quadrature of the cavity field, $c_{\mathrm{proj}}=\mathrm{Re}(\langle\hat{c}\rangle)\cos\theta+\mathrm{Im}(\langle\hat{c}\rangle)\sin\theta$, where $\theta=\arctan(-\kappa/\mathrm{\langle\widetilde{\Delta}_{pc}\rangle})$, defined with respect to the phase of the pump field. One finds
$c_{\mathrm{proj}}$ $=\frac{N\mathrm{\Omega}g_{0}}{2|\mathrm{\Delta_{pa}}|}\Bigg{\langle}\frac{\mathrm{\hat{\Theta}}}{\sqrt{\mathrm{\widetilde{\Delta}_{pc}}^{2}+\kappa^{2}}}\Bigg{\rangle}.$ (1)
We define $\mathrm{\widetilde{\Delta}_{pc}}=\mathrm{\Delta_{pc}}-\sum_{n=1}^{N}g_{0}^{2}\sin^{2}(k\hat{z}_{n})/\mathrm{\Delta_{pa}}$ as the effective cavity detuning including the dispersive shift from the atoms. The cavity field is proportional to the operator
$\hat{\Theta}\equiv\frac{1}{N}\sum_{n=1}^{N}\sin(k\hat{z}_{n}),$ (2)
whose mean value can be interpreted as the order parameter. For atoms trapped at the nodes $\langle\hat{\Theta}\rangle=0$, while for atoms at the antinodes $\langle\hat{\Theta}\rangle=\pm 1$. Thus $\langle\hat{\Theta}\rangle$ takes the role of an effective magnetization for the system [26, 42].
To detect the complex cavity field $\hat{c}$, we send $y$-polarized emission from one cavity mirror to a heterodyne receiver (see fig. S1) and record the signal for a total time of $250\,\mu\mathrm{s}$, including the $50\,\mu$s pump ramp-up time. We filter the time trace with an adjustable averaging time, ranging from $5\,\mu\mathrm{s}$ to $50\,\mu\mathrm{s}$; we average for a minimum of $5\,\mu$s because increased noise at shorter averaging times compromises the signal-to-noise ratio of the extracted cavity field. The values obtained at each step of the filter are treated as individual data points for $\hat{c}$. We determine the angle $\theta$ by referencing the phase of the distribution of $\hat{c}$ to that obtained with all atoms placed at antinodes of the same sign. We then project each value of $\hat{c}$ onto the axis defined by $\theta$ to obtain $c_{\mathrm{proj}}$ (see materials and methods and fig. S2). Our measured $c_{\mathrm{proj}}$ is scaled such that $c_{\mathrm{proj}}^{2}$ is the average photon number in the cavity.
To observe bifurcation in the distributions of $c_{\mathrm{proj}}$, we repeat the measurement many times at different pump strengths and analyze the heterodyne data from $t=0$ to $t=100\,\mu$s (where $t=0$ is the end of the pump ramp-up) using a $5\,\mu$s averaging time. Due to heating of the atoms from continuous exposure to pump light [43, 44], we constrain this analysis to the $100\,\mu$s span described above. The distributions of $c_{\mathrm{proj}}$ show a striking bifurcation as the pump strength is increased. An example is shown in Fig. 1B. We see that once $\mathrm{\Omega}$ surpasses a threshold near $2\pi\times 30\,\mathrm{MHz}$, the single peak at $c_{\mathrm{proj}}=0$ splits into two symmetric branches with finite field amplitude.
The simplest model that captures this physics is purely mechanical and lossless, with an effective Hamiltonian
$\begin{split}\hat{H}_{\text{eff}}\equiv&\sum_{n=1}^{N}\left[{\frac{\hat{p}_{n}^{2}}{2M}+\frac{M}{2}\nu^{2}\left({\hat{z}_{n}-z_{0n}}\right)^{2}}\right]+\hbar\mathcal{D}N^{2}\hat{\Theta}^{2}.\end{split}$ (3)
Here, $M$ is the atomic mass of 87Rb, $\nu=2\pi\times 93\,\mathrm{kHz}$ is the tweezer trap frequency (in the harmonic approximation), $\hbar$ is the reduced Planck constant, and $\mathcal{D}\equiv\left({\frac{g_{0}\Omega}{2\Delta_{\text{pa}}}}\right)^{2}\tilde{\Delta}_{\text{pc}}/(\tilde{\Delta}_{\text{pc}}^{2}+\kappa^{2})$ is the strength of the cavity-induced mechanical potential, whose sign depends on $\tilde{\Delta}_{\text{pc}}$. The tweezer positions are $z_{0n}=4.5\lambda\times n$.
However, thermal fluctuations as well as motion of the atoms perpendicular to the cavity axis are relevant for a quantitative description of the experiment. By means of a more refined optomechanical model [45], the critical pump strength is found to be
$\mathrm{\Omega_{c}}(T)=\sqrt{\frac{2M\nu^{2}\mathrm{\Delta^{2}_{pa}}(\mathrm{\Delta^{2}_{pc}}(T)+\kappa^{2})}{Ng^{2}_{0}k^{2}|\mathrm{\Delta_{pc}}(T)|\hbar\varepsilon(T)}},$ (4)
where $T$ is the atomic motional temperature and the temperature-dependent factor $\varepsilon(T)$ is defined in the supplementary text. The dependencies of $\mathrm{\Omega_{c}}$ on $N$ and $\mathrm{\Delta_{pa}}$ are studied in this work (see Fig. 2 and fig. S4). Equation (4) includes two different thermal effects on the critical pump strength. First, taking an average over the thermal motional shift of the atoms yields a net shift of the cavity resonance $\Delta_{\text{pc}}(T)\equiv\langle\tilde{\Delta}_{\text{pc}}\rangle$. Second, thermal fluctuations allow the atoms to probe the non-linearity of the effective mechanical potential in $\hat{\Theta}^{2}$. This leads to an increase in the critical pump strength proportional to the factor $1/\varepsilon(T)$ in Equation (4). Using a temperature of $35\pm 10\,\mu$K, determined from an independent measurement [38], Equation (4) predicts $\mathrm{\Omega_{c}}=2\pi\times(26.5\pm 2.5)\,$MHz for the parameter settings in Fig. 1B.
To determine $\mathrm{\Omega_{c}}$ quantitatively from the data, we draw inspiration from Landau’s theory of phase transitions [46]. If we only consider the contribution of the dominant mode to the Hamiltonian in Equation (3), the potential energy becomes $\frac{1}{2}NM\nu^{2}\hat{z}_{\mathrm{dom}}^{2}+\hbar\mathcal{D}N^{2}\sin^{2}(k\hat{z}_{\mathrm{dom}})$, where $\hat{z}_{\mathrm{dom}}$ is the displacement of the dominant mode with respect to the cavity nodes. If we consider the distribution of positions to be a Boltzmann distribution determined by the local potential energy, and expand the potential energy to fourth order in $\hat{z}_{\mathrm{dom}}$, we obtain a probability distribution $p(z_{\mathrm{dom}})$ of the form
$p(z_{\mathrm{dom}})\propto\exp(-Bz_{\mathrm{dom}}^{2}-Dz_{\mathrm{dom}}^{4}),$ (5)
where $B$ and $D$ are parameter-dependent constants and $z_{\mathrm{dom}}\equiv\langle\hat{z}_{\mathrm{dom}}\rangle$. According to Landau’s theory, a second order phase transition occurs at $B=0$. Because $c_{\mathrm{proj}}$ is proportional to $z_{\mathrm{dom}}$ to leading order, we fit the observed distributions of $c_{\mathrm{proj}}$ to Equation (5) and extract the coefficient $B$ to determine where $\mathrm{\Omega_{c}}$ occurs. Analyzing the distributions of $c_{\mathrm{proj}}$, we observe that the fits to the Boltzmann distribution match the data well (see fig. S5 and supplementary text) even though the system is driven and dissipative. From the data in Fig. 1B, we extract an experimental $\mathrm{\Omega_{c}}=2\pi\times(33.1^{+2.0_{\mathrm{syst}}}_{-2.8_{\mathrm{syst}}}\pm 1.6_{\mathrm{stat}})\,\mathrm{MHz}$. This measurement is slightly above the prediction of Equation (4) for our independent measurement of the initial temperature of the trapped atoms. The slight discrepancy may be caused by heating of the trapped atoms during the experiment, owing to light scattering. Nonetheless, when statistical and systematic errors (primarily resulting from uncertainty in our calibration of $\Omega$ [45]) are taken into consideration, the measurement falls within the error of the prediction of Equation (4) for our independent temperature measurement.
Figure 2: Dependence of the critical point on atom number. (A) Bifurcation data for $N=10$ (circles), $N=16$ (squares), and $N=22$ (triangles). Shaded areas show the width of the distribution at half maximum while markers show the maxima of the fitted Boltzmann distrbution. (B) Experimentally measured $\mathrm{\Omega_{c}}$ (circles) for different $N$ plotted on a logarithmic scale and the theoretical prediction of Equation (4), based on an independent temperature measurement of $T=35\pm 10\,\mu$K (black dashed line and light blue shaded area). Error bars represent statistical errors from fits of the parameter $B$ and drift of the pump power over the course of the experiment, estimated to be 10%. The reported values of $\Omega_{\mathrm{c}}$ account for systematic errors due to Gaussian detector noise and the Gaussian intensity profile of the pump beam. We note that there is an additional systematic error of +6.0% and -8.5% on our calibration of $\Omega$ (see supplementary text for details). All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$.
II Mesoscopic signatures of a phase transition
To probe the evolution of system properties with size directly, we examine how critical behavior varies with atom number. We first examine the dependence of $\mathrm{\Omega_{c}}$ on $N$, and find that $\Omega_{c}$ diminishes as $N$ increases. Quantitatively, for an averaging time of $5\,\mu\mathrm{s}$, we almost recover the $\sim 1/\sqrt{N}$ dependence predicted by Equation (4), shown in Fig. 2B. Though the critical point’s scaling with $N$ has been measured before in macroscopic bosonic and fermionic realizations of the superradiant phase transition [43, 47], our work shows that the $N$-dependence is also observable in a regime far from the thermodynamic limit. In Fig. 2B, the data appear above the prediction of Equation (4) if we use the temperature from our independent measurement, implying that the atoms have been heated to a temperature close to $66\,\mu$K at the time of observation.
Figure 3: Lifetime of the self-organized state. Above the critical point, we observe three dynamical behaviors in the cavity field time traces: (A) symmetry breaking, (B) switching, and (C) rapid oscillations. Time is measured from the end of the pump ramp-up. Shaded areas show the time spent in the $c_{\mathrm{proj}}>0$ state (red) and the $c_{\mathrm{proj}}<0$ state (blue). Shaded purple regions indicate the shot noise level. Data are shown for $N=20,\,\mathrm{\Omega}=2\pi\times 45.4\,\mathrm{MHz}$. (D) $g^{(1)}(0,t)$ for $N=20$ at $\mathrm{\Omega}<\mathrm{\Omega_{c}}$ and $\mathrm{\Omega}>\mathrm{\Omega_{c}}$. When the system is self-organized, $g^{(1)}(0,t)$ decays on a much longer timescale. All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$. (E) Dependence of the critical point on $N$ for different averaging times. At larger $N$, the curves converge, indicating that self-organization persists on longer timescales. At lower $N$, the separation between the curves grows due to the transience of the mesoscopic system. The measurements at $25\,\mu$s and $50\,\mu$s for $N<14$ are absent because the transition was not reached in the range of $\mathrm{\Omega}$ values scanned. The plotted $\Omega_{\mathrm{c}}$ data are subject to the same systematic errors as described in Fig. 2B. Lines are guides to the eye.
Atomic self-organization in a cavity can be compared with the crossover between a magnetized and unmagnetized state in a spin system. Notably, due to the infinite-range character of the interaction, the crossover does not occur by forming domains of increasing size as the critical coupling $\Omega_{\mathrm{c}}$ is approached [42]. Instead, the behavior of $\langle\hat{\Theta}\rangle$ above the transition is analogous to that of a superparamagnet: As discussed in seminal work by Néel [48], small ferromagnetic particles are subject to thermal fluctuations that cause their magnetization to relax on a time scale that grows with the particle volume. Averaging the magnetization over times much longer than the relaxation time results in zero net magnetization, while averaging over times much shorter than the relaxation time allows for observation of quasistatic magnetic ordering [5]. We similarly find that mesoscopic self-organization of atoms has a characteristic lifetime that depends on system size.
To reveal the dynamics of the magnetization order parameter $\langle\hat{\mathrm{\Theta}}\rangle$, we take advantage of the cavity’s $\mu$s-scale readout capability [35]. In Fig. 3, we examine single-shot time traces of $c_{\mathrm{proj}}$, time correlations of the field, as well as the $N$-dependence of the critical point as we increase the averaging time. Fig. 3A-C show single-shot time traces representative of observed dynamical behaviors in the self-organized regime: symmetry breaking, switching, and rapid oscillations. For a macroscopic atom number, only symmetry breaking should occur, while for a mesoscopic atom number, fluctuations away from equilibrium can lead to switching between stable states of the system [42]. The rapid oscillation events are rare, occurring in only 2 out of the 180 shots obtained at the shown parameter setting; we expect they result from the energy in the dominant mode being above the energy gap separating the two symmetry-broken states.
To highlight the difference in dynamical behavior below and above the critical point, we calculate the first-order coherence function of the field quadrature $g^{(1)}(t_{1},t_{2})$, defined as $g^{(1)}(t_{1},t_{2})=\langle c^{*}_{\mathrm{proj}}(t_{1})c_{\mathrm{proj}}(t_{2})\rangle/\sqrt{\langle|c_{\mathrm{proj}}(t_{1})|^{2}\rangle\langle|c_{\mathrm{proj}}(t_{2})|^{2}\rangle}$, where the angle brackets denote averaging over multiple shots. In particular, $g^{(1)}(0,t)$ indicates how correlated the field at time $t_{2}=t$ is to the field at time $t_{1}=0$, where $t_{1}=0$ is chosen to be at the end of the pump ramp-up. Symmetry breaking is indicated by $g^{(1)}(0,t)$ remaining high for extended $t$. In Fig. 3D, we see that when $\mathrm{\Omega}/\mathrm{\Omega_{c}}=0.8$, $g^{(1)}(0,t)$ decays rapidly, indicating the absence of symmetry breaking, and exhibits oscillations at approximately $60\,\mathrm{kHz}$, corresponding to the softened trap frequency. At $\mathrm{\Omega}/\mathrm{\Omega_{c}}=1.6$, $g^{(1)}(0,t)$ decays on a much longer timescale and oscillations are absent, indicating a prolonged lifetime of the symmetry-broken states.
We next observe the $N$-dependence of the lifetime of the self-organized state. As discussed above, the characteristic behavior of the system depends on the time over which $\langle\hat{\mathrm{\Theta}}\rangle$ is averaged. On short time scales, the system displays signatures of symmetry breaking, while on long time scales, the system switches between the two stable states. As a result, in Fig. 3E, we see that the nominal $\mathrm{\Omega_{c}}$ increases as we increase the averaging time from 5 to $50\,\mu\mathrm{s}$. Remarkably, we see that $\mathrm{\Omega_{c}}$ increases significantly at lower $N$ but less so at higher $N$. This indicates that as the system size decreases, the characteristic time scale over which the system remains self-organized also decreases. We thus observe a key feature of this mesoscopic system: Averaging $\langle\hat{\mathrm{\Theta}}\rangle$ over times longer than the $N$-dependent lifetime erases signatures of symmetry breaking.
Figure 4: Optomechanical susceptibility. (A) Schematic of the susceptibility measurement. The dominant mode of the array ($z_{\mathrm{dom}}$) is displaced by $\delta z$. When $\delta z=0$, the effective potential $V_{\mathrm{eff}}$ on the atoms is symmetric about $z_{\mathrm{dom}}=0$ (i). For $|\delta z|>0$, $V_{\mathrm{eff}}$ tilts in the direction of the bias (ii-iii). (B) Distributions of $c_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ (small purple points) and their mean values (circles) at various $\delta z$ for $N=18,\,\mathrm{\Omega/\Omega_{c}}=1.16$. Measurement of $\partial(\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}})/\partial(\delta z)\big{\rvert}_{\delta z=0}$ is taken by fitting the slope of $\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ for $-0.01\lambda\leq\delta z\leq 0.01\lambda$ (red circles). (C) Susceptibility $\chi$ at various $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ for $N=14$ and $N=18$. The data (hexagons and circles) are plotted together with the theoretical prediction based on the Boltzmann distribution (dashed lines). All data in figure are taken at $\mathrm{\Delta_{pa}}=-2\pi\times 80\,\mathrm{MHz}$ and $\mathrm{\Delta_{pc}}=-2\pi\times 1.9\,\mathrm{MHz}$.
III Finite susceptibility at criticality
The positional control afforded by the tweezers not only allows us to observe spontaneous symmetry breaking, but also to explicitly break symmetry with nanometer-scale precision and probe the system’s response. By displacing adjacent tweezers in opposite directions by an amount $\delta z$, we introduce a constant force proportional to $\delta z$ on the atoms that biases the dominant mode (Fig. 4A). Generally, the amount a system moves in response to an applied force is characterized by its mechanical susceptibility. In our case, $c_{\mathrm{proj}}$ serves as a proxy for the atom positions. As such, we obtain the optomechanical susceptibility, $\chi$, which we define as
$\chi\equiv\frac{1}{k\bar{c}_{\mathrm{a.n.}}}\frac{\partial\bar{c}_{\mathrm{proj}}}{\partial(\delta z)}\bigg{\rvert}_{\delta z=0}.$ (6)
Here, $\bar{c}_{\mathrm{a.n.}}$ is the measurement of $\bar{c}_{\mathrm{proj}}$ when the tweezers are positioned at the positive antinodes. The overbar indicates an average over the distribution of $c_{\mathrm{proj}}$ obtained from the first $50\,\mu\mathrm{s}$ of multiple shots. The quantity $\partial(\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}})/\partial(\delta z)\big{\rvert}_{\delta z=0}$ is measured by fitting the slope of $\bar{c}_{\mathrm{proj}}/\bar{c}_{\mathrm{a.n.}}$ in the range $-0.01\lambda\leq\delta z\leq 0.01\lambda$ for $N=18$ (red circles in Fig. 4B) and $N=14$. We note that the atoms are initialized with root-mean-square thermal fluctuations in position of about 100 nm, leading to a thermal Lamb-Dicke parameter of about 0.8 [45].
From measurement of $\chi$ for $N=14$ and $N=18$ at different values of $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ (Fig. 4C), we make two observations: 1) while $\chi$ increases with $\mathrm{\Omega/\Omega_{c}}$, it does not diverge at the critical point as one would expect for a second-order phase transition in a macrosopic system, and 2) the value of $\chi$ grows with atom number for $\mathrm{\Omega/\Omega_{c}}>1$. These observations can be explained by a model assuming $\chi\simeq\partial\langle z_{\mathrm{dom}}\rangle/\partial(\delta z)\big{\rvert}_{\delta z=0}=NM\nu^{2}\langle\Delta z_{\mathrm{dom}}^{2}\rangle/(k_{B}T)\big{\rvert}_{\delta z=0}$, where angle brackets indicate a thermal average including the tweezer bias $\delta z$[45]. Thus, the susceptibility is directly related to the size of fluctuations in the system through the variance $\langle\mathrm{\Delta}z_{\mathrm{dom}}^{2}\rangle$. In a macroscopic system, one would expect $\chi$ to diverge according to $1/(1-(\mathrm{\Omega/\Omega_{c}})^{2})$ as $\mathrm{\Omega}$ approaches $\mathrm{\Omega_{c}}$. Instead, because of our finite atom number, our model predicts that $\chi$ should remain finite at the critical point.
Specifically, at the critical point, our model predicts that $\chi$ should be approximately proportional to $\sqrt{N}$ as a result of the variance $\langle\mathrm{\Delta}z_{\mathrm{dom}}^{2}\rangle$ scaling as $1/\sqrt{N}$ at leading order. This has a simple interpretation: At smaller $N$, the fluctuations in $z_{\mathrm{dom}}$ are larger and the atoms sample the quartic part of the effective optomechanical potential $V_{\mathrm{eff}}$, which provides the confinement that prevents the susceptibility from diverging. While we observe finite $\chi$ at the critical point, we are unable to resolve the $\sqrt{N}$-dependence due to the quality of our data. Nonetheless, we emphasize that this observation of finite $\chi$ is a direct measure of mesoscopic fluctuations and not just a result of non-zero temperature—a macroscopic system at non-zero temperature would still exhibit a diverging susceptibility. As $\mathrm{\Omega}/\mathrm{\Omega_{c}}$ is increased beyond 1, our model predicts that $\chi$ should continue to grow and eventually scale proportionally to $N$, before saturating at a value of $NM\nu^{2}\lambda^{2}/(16k_{B}T)$. For $\mathrm{\Omega}\gg\mathrm{\Omega_{c}}$, the $\delta z$ required to confine $\langle z_{\mathrm{dom}}\rangle$ effectively to one of the minima of $V_{\mathrm{eff}}$ scales inversely with $N$, leading to $\chi\propto N$. We note that $\chi$ continuing to increase for $\Omega>\Omega_{\mathrm{c}}$ is a consequence of us applying the tweezer shift $\delta z$ to break symmetry before the cavity interaction is turned on, analogous to applying a magnetic field to a spin system before lowering its temperature below the Curie point. In contrast, if the tweezer shift were applied after the cavity interaction, one might expect $\chi$ to peak at $\Omega=\Omega_{\mathrm{c}}$. Accordingly, we see the separation between the $N=14$ and $N=18$ data become more pronounced as we increase $\mathrm{\Omega}$ beyond $\mathrm{\Omega_{c}}$, consistent with the prediction of $\chi$ approaching linear scaling with $N$. These results highlight how the response of mesoscopic systems deviates from that of macroscopic systems.
IV Discussion and outlook
We have studied self-organization, bifurcation, and symmetry breaking as well as mesoscopic dynamics and response due to light-mediated interactions in an atom array. The deterministic atom number and precise control over interactions necessary for this work were achieved by interfacing an optical tweezer array with an optical cavity, opening new directions for these systems. The realization of mesoscopic self-organization with control over individual atoms has significant implications not just for our fundamental understanding of mesoscopic systems, but also for quantum simulation, optimization, and metrology. On one hand, the dynamics of the cavity field provide a plethora of information that may further our understanding of nonequilibrium physics in mesoscopic systems [49] and pave the way for time-resolved quantum simulation [50, 51]. On the other hand, mesoscopic self-organization may also open the door to programmable engineering [52] of exotic symmetry-breaking phases of matter [53], for instance by illuminating groups of atoms with different phases of pump light. This system can also be used to implement a quantum numerical solver for a variety of classically hard computational problems [54, 55, 56], where the tweezer array allows for highly programmable initial conditions. Additionally, the optomechanical susceptibility near the critical point can be a resource for enhanced sensing of applied forces [57, 58, 59].
V Acknowledgments
We thank Nathan Song for assistance in the lab.
Funding: We acknowledge support from the AFOSR (Grant No. FA9550-1910328 and Young Investigator Prize Grant No. 21RT0751), from ARO through the MURI program (Grant No. W911NF-20-1-0136), from DARPA (Grant No. W911NF2010090), from the NSF (QLCI program through grant number OMA-2016245, and CAREER Award No. 2047380), from the David and Lucile Packard Foundation, and from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. J.H. acknowledges support from the Department of Defense through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program. C.C.R. acknowledges support from the European Union’s Horizon Europe program under the Marie Sklodowska Curie Action LIME (Grant No. 101105916).
Author contributions: J.H., Y.-H.L., Z.Y., and T.X. contributed to building the experimental setup, performing the experiments, and analyzing the data. C.C.R., S.J.M., A.A.-G., D.M.S.-K., and J.H. contributed to the theoretical model. Z.Y. and D.M.S.-K. conceived the experiments. All authors contributed to writing of the manuscript and discussed the results. All work was supervised by D.M.S.-K. and A.A.-G.
Competing interests: The authors declare no competing interests.
Data and materials availability: Data are available from the corresponding author upon reasonable request.